Find the intercepts of -x-y=2 and graph the line.
Answer by mathgranny(7)

Find the intercepts of -x-y=2 and graph the line.
The x-intercept is on the x-axis, the point where the line crosses the x-axis.
Plot a few points on the x-axis and list the coordinates of the points.
Which coordinate is always 0?
For points on the x-axis, the second coordinate is 0, so Y = 0.
In the above equation set y = 0, and solve for x.
-x - y = 2.
-x - 0 = 2
-X = 2
S0 +X = -2
(-2, 0) Is the x-intercept which is on the line.

TO FIND THE Y - INTERCEPT, SET X = 0
-(0) - Y = 2
-Y = 2
SO +Y = -2
AND THE POINT (0, -2) is on the line.
Plot these two points and draw a line through them.
You can also find the slope by calculating the rise over the run.
Remember that each point that is the intercept has 0 for one coordinate. students sometimes get this backwards.
Just remember y-intercept? x = 0
x-intercept ? y = 0